Please answer questions a b and c
I0 a) If a fair die is rolled 22 times, what is the probability that a 6 is obtained exactly 3 times? (b) If a fair die is rolled, what is the probability that the third time that a 6 is obtained is on the tenth roll? (c) If a fair coin is tossed 11 times, what is the prob ability that three or fewer heads are obtained?
Answer #1

Please note nCx = n! / [(n-x)!*x!]

(a) Binomial Probability = nCx * (p)x * (q)n-x, where n = number of trials and x is the number of successes.

Here n = 22, x = 3, p (Probability of getting a 6) = 1/6 and q = 1 - p = 1 - 1/6 = 5/6

P(3) = 22C3 * (1/6)3 * (5/6)22-3 = 19 = 0.2232

(b) Here since we are looking for the 3rd success (which is > 1 success). We use the negative Binomial Distribution

P[X = x] = x-1 C k-1 * (p)k * (q) x-k, where k is the success value, x is the number of trials.

Here x = 10 and k = 3. p = 1/6 and q = 5/6

Therefore P [X = 10] = 10-1 C 3-1 x (1/6)3 x (5/6) 10-3 = 7 = 0.0465